### v1

# On point spectrum of Jacobi matrices generated by iterations of quadratic polynomials

Abstract.
In general, point spectrum of an almost periodic Jacobi matrix can depend on the element of the hull. In this paper, we study the hull of the limit-periodic Jacobi matrix corresponding to the equilibrium measure of the Julia set of the polynomial $z^{2}-\lambda$ with large enough $\lambda$; this is the leading model in inverse spectral theory of ergodic operators with zero measure spectrum. We prove that every element of the hull has empty point spectrum.
Key words and phrases:
Jacobi matrix, point spectrum, limit periodic operator, hull, Ruelle operator
1991 Mathematics Subject Classification:
47B36, 42C05, 37C30, 37F15
B. E. was supported by the Austrian Science Fund FWF, project no: P33885
M. L. was supported in part by NSF grant DMS–2154563.
P. Y. was supported by the Austrian Science Fund FWF, project no: P34414
1. Introduction
In an ergodic family of operators, ergodicity implies almost sure constancy of the spectrum, as well as a.c., s.c., p.p. spectrum [17]. Historically, the a.c. spectrum was first understood in the setting of periodic operators, and the pure point spectrum gained prominence in the context of Anderson localization. Singular continuous spectrum was originally perceived as an anomaly, but then discovered to be generic in many settings [20, 3, 8]. The study of spectral type of ergodic operators is now a vast area of research; it allows only a few general results, such as Kotani theory [15, 19], and a lot of research is centered on prominent models such as the Anderson model, the Almost Mathieu operator, and the Fibonacci Hamiltonian; the dependence of spectral type on the ergodic space and the size and arithmetic properties of the parameters are of interest, and the analyses are highly model-dependent.
Let us focus on the setting of almost periodic Jacobi matrices, recalling that a Jacobi matrix
$(Ju)_{n}=a_{n}u_{n-1}+b_{n}u_{n}+a_{n+1}u_{n+1}$
is almost periodic if the set of its translates $S^{n}JS^{-n}$ is precompact in the set of bounded operators on $\ell^{2}({\mathbb{Z}})$; here $S$ denotes the shift on $\ell^{2}({\mathbb{Z}})$. The closure of the set of translates is called the hull of $J$. Equivalently, $J$ is almost periodic if there exist continuous maps ${\mathcal{A}},{\mathcal{B}}:G\to{\mathbb{R}}$ on a compact Abelian group $G$ with ${\mathcal{A}}\geq 0$ and $\omega,\theta\in G$ such that
$a_{n}={\mathcal{A}}(\theta+n\omega),\qquad b_{n}={\mathcal{B}}(\theta+n\omega).$
By passing to a subgroup, one reduces to the case when the orbit $\{\theta+n\omega\mid n\in{\mathbb{Z}}\}$ is dense in $G$ and obtains an ergodic family, with the hull in bijective correspondence to the phase $\theta\in G$. For almost periodic operators, some of the ”almost sure” statements extend to ”sure” statements: the spectrum is constant, and by a result of Last–Simon [16], the absolutely continuous spectrum is constant. However, other spectral types are more sensitive to perturbations, and can be altered by changing the phase $\theta$.
For instance, the Almost Mathieu operator
$(Ju)_{n}=u_{n-1}+2\lambda\cos(\theta+n\omega)u_{n}+u_{n+1},$
with irrational $\omega\in{\mathbb{T}}$ is an almost periodic operator, with hull indexed by $\theta\in{\mathbb{T}}$. In the subcritical case $\lambda<1$, it has a.c. spectrum, but in the supercritical case $\lambda>1$, its spectrum is s.c. or p.p. depending on the phase [14, 12]. A remarkable result of Jitomirskaya [13] proves that the critical ($\lambda=1$) Almost Mathieu operator has empty p.p. spectrum for every choice of $\theta$, after partial results [2].
Jacobi matrices generated by iterations of quadratic polynomials are the leading model in inverse spectral theory. Starting from the pioneering work of Bellisard–Bessis–Moussa [6], they have attracted much attention, providing almost periodic operators with singular continuous spectrum; cf. [4, 5, 7, 18]. In this paper, we prove absence of pure point spectrum in the hull for this model and view this as a counterpart of Jitomirskaya’s theorem.
To describe the model, start with an even quadratic polynomial $T(z)=z^{2}-\lambda$, $\lambda>2$. Its Julia set
${\mathsf{E}}={\mathbb{C}}\setminus\{z\in{\mathbb{C}}\mid T^{\circ n}(z)\to\infty,n\to\infty\}$
is a subset of ${\mathbb{R}}$ of zero one-dimensional Lebesgue measure; we denote $T^{\circ n}=T\circ T^{\circ(n-1)}$. Denote by $\mu_{\mathsf{E}}$ the equilibrium measure for the set ${\mathsf{E}}$. The measure $\mu_{\mathsf{E}}$ generates orthogonal polynomials which obey the Jacobi recurrence relation
$xp_{n}(x)=a_{n}p_{n-1}(x)+a_{n+1}p_{n+1}(x)$
(note that $b_{n}=0$ by symmetry), with $a_{0}=0$. For $\lambda>3$, the coefficients $a_{n}$ form a limit periodic sequence: for an arbitrary $\epsilon>0$ there exist $n_{0}$ such that
$|a_{k+s2^{n}}-a_{k}|\leq\epsilon,\quad\forall n\geq n_{0}$
uniformly in $k$ and $s$. Due to this property the sequence can be extended to a continuous function on dyadic integers ${\mathbb{Z}}_{2}$
$a_{\varkappa}=\lim_{k_{j}\to{\varkappa}}a_{k_{j}},\quad k_{j}\to{\varkappa}\in{\mathbb{Z}}_{2}.$ (1.1)
This extends the half-line Jacobi matrix associated to $\mu_{\mathsf{E}}$ to a full-line limit-periodic Jacobi matrix $J_{0}$, and describes the hull of this matrix as $\{J_{\varkappa}\mid{\varkappa}\in{\mathbb{Z}}_{2}\}$ where
$J_{\varkappa}e_{n}=a_{{\varkappa}+n}e_{n-1}+a_{{\varkappa}+n+1}e_{n+1},$
and $\{e_{n}\}_{n\in{\mathbb{Z}}}$ is the standard basis in $\ell^{2}({\mathbb{Z}})$.
All $J_{\varkappa}$ have spectrum ${\mathsf{E}}$, so they have purely singular spectrum. It was proved [21] that $J_{0}$ has no eigenvalues; this proof used the observation $a_{0}=0$ to decompose $J_{0}$ as a direct sum of half-line Jacobi matrices, and used Ruelle theory to study the other half-line matrix. Since $a_{\varkappa}>0$ for ${\varkappa}\neq 0$, $J_{\varkappa}$ has no such decomposition if ${\varkappa}\notin{\mathbb{Z}}$. Nonetheless, we prove the following:
Theorem 1.1.
For $\lambda>3$ and $\varkappa\in{\mathbb{Z}}_{2}$, the operator $J_{\varkappa}$ on $\ell^{2}({\mathbb{Z}})$ has no eigenvalues. It has purely singular continuous spectrum.
The key tool are matrix Ruelle operators, see Definition 2.3. A matrix generalization of Ruelle theory [9] seems a hard problem, due to noncommutativity. For instance, whereas the multiplicative representation of Schur functions, after separation of a Blaschke product, is a simple observation, its matrix valued analog is the multiplicative theory of $J$-contractive matrices [1]. Note that de Branges theory of canonical systems [10] is its partial case. However, we were able to reduce our problem to matrix Ruelle operators with special rational weights, and due to this, our construction becomes parallel to that suggested in [21].
It is an open problem whether the spectral measure corresponding to an element $\varkappa$ with periodic digits, such as $\varkappa=1+\sum_{j=0}^{\infty}2^{2j}$ (in ${\mathbb{Z}}_{2}$), is mutually singular to the spectral measure of $J_{0}$. The operator $J_{\varkappa}$ corresponds to a matrix Ruelle operator, and using Ruelle theory for analogy suggests that different weights should correspond to mutually singular measures. If so, this would be another example of operators in the same hull which are not unitarily equivalent, in strong contrast to the results of [22] for reflectionless operators on regular Widom spectra with DCT property, where all operators in the same hull are unitarily equivalent to each other.
Besides the intrinsic interest in the spectral type, absence of eigenvalues is used in an upcoming preprint [11] in the analysis of the scaling behavior of Christoffel–Darboux kernels corresponding to $\mu_{\mathsf{E}}$.
Some explicit formulas are omitted from the journal version of this paper, and are included in the arXiv version.
We thank David Damanik for helpful information.
2. Hull $\{J_{\varkappa}\}$ and matrix Ruelle operators
2.1. Background: Ruelle operators and half-line Jacobi matrices
To a Hölder function $\varphi$ one associates the Ruelle operator
$({\mathcal{L}}_{\varphi}f)(x)=\sum_{T(y)=x}e^{{\varphi}(y)}f(y),\quad f\in C({\mathsf{E}}).$
Let $\rho=\rho(\varphi)$ be the spectral radius of ${\mathcal{L}}_{\varphi}$. According to the Ruelle version of the Perron-Frobenius theory, $\rho$ is an eigenvalue for the operator and its adjoint
$\frac{1}{\rho}{\mathcal{L}}_{\varphi}h=h,\quad\frac{1}{\rho}{\mathcal{L}}^{*}_{\varphi}\nu=\nu,$
with a positive continuous function $h\in C({\mathsf{E}})$ and a (positive) measure $\nu$. In particular, the equilibrium measure $\mu_{\mathsf{E}}$ for the set ${\mathsf{E}}$ is an eigenvector for ${\mathcal{L}}_{0}^{*}$,
$\int f(x)\mu_{\mathsf{E}}(dx)=\frac{1}{2}\int({\mathcal{L}}_{0}f)(x)\mu_{\mathsf{E}}(dx).$ (2.1)
We denote by $J_{+}$ the half-line Jacobi matrix corresponding to $\mu_{\mathsf{E}}$.
The extension of $J_{+}$ to a full-line operator $J_{0}$ uses the continuity relation (1.1). In particular, the extension of the sequence $(a_{k})_{k\in{\mathbb{Z}}_{+}}$ to $(a_{k})_{k\in{\mathbb{Z}}}$ is given by
$a_{-k}=\lim_{n\to\infty}a_{2^{n}-k}.$
Since $a_{0}=0$, this gives a direct sum decomposition $J_{0}=J_{+}\oplus J_{-}$. We use +=${\mathbb{Z}}_{+}=\{n\in{\mathbb{Z}}\mid n\geq 0\}$ and ${\mathbb{Z}}_{-}={\mathbb{Z}}\setminus{\mathbb{Z}}_{+}$.
It was proved in [21] that the canonical spectral measure $\nu(dx)$ of $J_{-}$ is the eigenmeasure of the adjoint to the Ruelle operator
${\mathcal{L}}_{2}f(x)=\sum_{T(y)=x}\frac{f(y)}{T^{\prime}(y)^{2}},$
i.e., ${\mathcal{L}}_{2}^{*}\nu=\rho_{2}\nu$ and $\rho_{2}=\frac{1}{2a_{-1}^{2}}$. Related orthogonal polynomials $P_{n}(x,\nu)$ possess the renorm relation
$P_{2n+1}(x,\nu)=\frac{x}{a_{-1}}P_{n}(T(x),\nu).$ (2.2)
2.2. Shift in ${\mathbb{Z}}_{2}$ and associated to $J_{\varkappa}$ matrix Ruelle operators
Lemma 2.1.
Let $Ve_{n}=e_{2n}$. The following identity holds
$V^{*}(J_{2{\varkappa}}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{\varkappa}-T(z))^{-1}$ (2.3)
Proof.
Due to (2.1) the operator $f(x)\mapsto f(T(x))$ is an isometry in $L^{2}_{\mu_{{\mathsf{E}}}}$. Its adjoint is $\frac{1}{2}{\mathcal{L}}_{0}$ viewed as an operator in $L^{2}_{\mu_{{\mathsf{E}}}}$. With the standard eigenfunction expansion $\ell^{2}({\mathbb{Z}}_{+})\to L^{2}_{\mu_{\mathsf{E}}}$, $e_{n}\mapsto p_{n}(\cdot,\mu)$, the operator $f\mapsto f\circ T$ is conjugated to the operator $V$ on $\ell^{2}({\mathbb{Z}}_{+})$. Thus the identity in the spectral representation
$\frac{1}{2}\sum_{T(y)=x}\frac{f(T(y))}{y-z}=\frac{1}{2}\frac{T^{\prime}(z)}{x-T(z)}f(x)$
can be rewritten as
$V^{*}(J_{+}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{+}-T(z))^{-1}.$
Likewise, the standard eigenfunction expansion for $J_{-}$ can be written as $\ell^{2}({\mathbb{Z}}_{-})\to L^{2}_{\nu}$, $e_{-1-n}\mapsto p_{n}(\cdot,\nu)$, and due to (2.2), the adjoint of $V$ on $\ell^{2}({\mathbb{Z}}_{-})$ in the spectral representation takes the form
$a_{-1}{\mathcal{L}}_{1}f(x)=a_{-1}\sum_{T(y)=x}\frac{f(y)}{T^{\prime}(y)},\quad f\in L^{2}_{\nu}.$
Similarly to $J_{+}$, the identity
$a_{-1}\sum_{T(y)=x}\frac{1}{y-z}\frac{yf(T(y))}{a_{-1}T^{\prime}(y)}=\frac{1}{2}\sum_{T(y)=x}\frac{1}{y-z}f(x)=\frac{1}{2}\frac{T^{\prime}(z)}{x-T(z)}f(x)$
yields the identity
$V^{*}(J_{-}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{-}-T(z))^{-1}.$
Taking direct sums yields
$V^{*}(J_{0}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{0}-T(z))^{-1}$
for operators on $\ell^{2}({\mathbb{Z}})$, and since $VS=S^{2}V$ we have
$V^{*}(S^{-2k}J_{0}S^{2k}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(S^{-k}J_{0}S^{k}-T(z))^{-1}.$
Therefore for a sequence $k_{s}\to{\varkappa}\in{\mathbb{Z}}_{2}$ we have $S^{-k_{s}}J_{0}S^{k_{s}}\to J_{\varkappa}$ and consequently (2.3). ∎
Let ${\mathcal{E}}:{\mathbb{C}}^{2}\to\ell^{2}({\mathbb{Z}})$ be defined by
${\mathcal{E}}\begin{pmatrix}C_{-1}\\ C_{0}\end{pmatrix}=\begin{pmatrix}e_{-1}&e_{0}\end{pmatrix}\begin{pmatrix}C_{-1}\\ C_{0}\end{pmatrix}=C_{-1}e_{-1}+C_{0}e_{0}.$
The spectral matrix-measure $\Sigma_{\varkappa}(dx)$ associated to $J_{\varkappa}$ is defined by
${\mathcal{E}}^{*}(J_{\varkappa}-z)^{-1}{\mathcal{E}}=\int\frac{\Sigma_{\varkappa}(dx)}{x-z}.$
We introduce the shift operation $\underline{s}:{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}$ as
$\underline{s}{\varkappa}=\sum_{k=0}^{\infty}{\epsilon}_{k+1}2^{k},\quad{\varkappa}=\sum_{k=0}^{\infty}{\epsilon}_{k}2^{k}$
Lemma 2.2.
Let
$\hat{\underline{s}}{\varkappa}=\begin{cases}\underline{s}{\varkappa},&{\varkappa}=2\underline{s}{\varkappa}\\ 1+\underline{s}{\varkappa},&{\varkappa}=1+2\underline{s}{\varkappa}\end{cases}$
Define
${\mathfrak{p}}_{\varkappa}(x)=\begin{pmatrix}a_{{\varkappa}-1}&0\\ a_{{\varkappa}}&1\end{pmatrix}\begin{pmatrix}2/{T^{\prime}(x)}&0\\ 0&1\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},$
and
${\mathfrak{p}}_{\varkappa}(x)=\begin{pmatrix}1&a_{{\varkappa}}\\ 0&a_{{\varkappa}+1}\end{pmatrix}\begin{pmatrix}1&0\\ 0&2/{T^{\prime}(x)}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.$
The following renorm equation holds
$\Sigma_{{\varkappa}}(dx)=\frac{1}{2}{\mathfrak{p}}_{{\varkappa}}^{*}(x){\mathcal{L}}_{0}^{*}\Sigma_{\hat{}\underline{s}{\varkappa}}(dx){\mathfrak{p}}_{{\varkappa}}(x).$ (2.4)
Proof.
Let ${\varkappa}=2\underline{s}{\varkappa}$, i.e., ${\epsilon}_{0}=0$. We will pass to spectral matrix-measures. Recall that in the spectral representation for two-sided Jacobi matrices, for $k\geq 0$, we have
$e_{k}\mapsto\begin{pmatrix}-a_{0}q^{+}_{k}\\ p^{+}_{k}\end{pmatrix},\quad e_{-1-k}\mapsto\begin{pmatrix}p^{-}_{k}\\ -a_{0}q^{-}_{k}\end{pmatrix},$
where $p_{k}^{\pm},q_{k}^{\pm}$ are orthogonal polynomials of the first and second type generated by $(J_{\varkappa})_{\pm}$. In particular
$e_{-2}\mapsto\begin{pmatrix}\frac{x}{a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix}=\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}-1}}\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}\end{pmatrix},\quad e_{0}\mapsto\begin{pmatrix}0\\ 1\end{pmatrix}.$
Therefore (2.3) implies
$\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}-1}}&0\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}&1\end{pmatrix}^{*}\Sigma_{\varkappa}(dx)\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}-1}}&0\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}-1}}&1\end{pmatrix}=\frac{1}{2}{\mathcal{L}}_{0}^{*}\Sigma_{\underline{s}{\varkappa}}(dx).$
We get (2.4) in the given case.
Let ${\varkappa}=1+2\underline{s}{\varkappa}$, i.e., ${\epsilon}_{0}=1$. Now we have
$V^{*}S(J_{{\varkappa}}-z)^{-1}S^{-1}V=V^{*}(J_{{\varkappa}-1}-z)^{-1}V=\frac{1}{2}T^{\prime}(z)(J_{\underline{s}{\varkappa}}-T(z))^{-1}.$ (2.5)
We define an involution $\tau e_{n}=e_{-1-n}$, $\tau^{2}=I,\tau^{*}=\tau$. Note that
$\tau V\tau e_{n}=\tau Ve_{-1-n}=\tau e_{-2-2n}=e_{2n+1}=SVe_{n}$
and
$\tau S\tau e_{n}=\tau Se_{-1-n}=\tau e_{-n}=e_{-1+n}=S^{-1}e_{n}.$
Denote $J^{\sharp}=\tau J\tau$. Conjugating (2.5) with $\tau$ we get
$V^{*}S^{-2}(J^{\sharp}_{{\varkappa}}-z)^{-1}S^{2}V=S^{-1}V^{*}(J^{\sharp}_{{\varkappa}}-z)^{-1}VS=\frac{1}{2}T^{\prime}(z)(J^{\sharp}_{\underline{s}{\varkappa}}-T(z))^{-1}$
and therefore (2.5) gets the form
$V^{*}(J^{\sharp}_{{\varkappa}}-z)^{-1}V=$ $\frac{1}{2}T^{\prime}(z)(SJ^{\sharp}_{\underline{s}{\varkappa}}S^{-1}-T(z))^{-1}$
$=$ $\frac{1}{2}T^{\prime}(z)(\tau S^{-1}J_{\underline{s}{\varkappa}}S\tau-T(z))^{-1}$
$=$ $\frac{1}{2}T^{\prime}(z)(J_{\underline{s}{\varkappa}+1}^{\sharp}-T(z))^{-1}.$ (2.6)
Using
$\tau{\mathcal{E}}={\mathcal{E}}\sigma_{1},\quad\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$
we have
$\Sigma^{\sharp}(dx)=\sigma_{1}\Sigma(dx)\sigma_{1},\quad\text{where}\quad\int\frac{\Sigma^{\sharp}(dx)}{x-z}={\mathcal{E}}^{*}(J^{\sharp}-z)^{-1}{\mathcal{E}}.$ (2.7)
Thus (2.2) yields the following relation for spectral matrix-functions
$\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}+1}}&0\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}+1}}&1\end{pmatrix}^{*}\Sigma_{\varkappa}^{\sharp}(dx)\begin{pmatrix}\frac{T^{\prime}(x)}{2a_{{\varkappa}+1}}&0\\ -\frac{a_{{\varkappa}}}{a_{{\varkappa}+1}}&1\end{pmatrix}=\frac{1}{2}{\mathcal{L}}_{0}^{*}\Sigma_{\underline{s}{\varkappa}+1}^{\sharp}(dx)$
Using (2.7), we obtain (2.4) in the second case. ∎
We will use duality with respect to the pairing
$\int\text{\rm tr}\{g(x)\Sigma(dx)\},$
where $g$ is $2\times 2$ continuous matrix function on ${\mathsf{E}}$. Based on this we define matrix Ruelle operators.
Definition 2.3.
We introduce the following Ruelle operator with matrix weight, acting on $2\times 2$ continuous matrix-functions
$({\mathcal{M}}_{\varkappa}g)(x)=\frac{1}{2}\sum_{T(y)=x}{\mathfrak{p}}_{{\varkappa}}(y)g(y){\mathfrak{p}}^{*}_{{\varkappa}}(y).$ (2.8)
Note that
$\int\text{\rm tr}\{g(x)({\mathcal{M}}_{\varkappa}^{*}\Sigma)(dx)\}=\int\text{\rm tr}\{({\mathcal{M}}_{\varkappa}g)(x)\Sigma(dx)\},$
and $g(x)\geq 0$ implies $({\mathcal{M}}_{\varkappa}g)(x)\geq 0$.
2.3. Action of Ruelle operators on a class of rational functions
In this subsection we show that the collection of rational functions with simple poles at $T^{\circ k}(0)$ is invariant for the action of ${\mathcal{M}}_{\varkappa}$ (at the moment we do not consider its closure).
Lemma 2.4.
Let
$g_{z}(x)=\frac{A}{x-z},\quad z\in{\mathbb{C}}\setminus{\mathsf{E}}$
with a constant $2\times 2$ matrix $A$. Then
$({\mathcal{M}}_{\varkappa}g_{T^{\circ k}(0)})(x)=\frac{\Phi_{\varkappa}A\Phi_{{\varkappa}}^{*}}{-T^{\circ k}(0)(x-T(0))}+\frac{T^{\circ k}(0){\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))A{\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))^{*}}{x-T^{\circ(k+1)}(0)},$ (2.9)
where
$\Phi_{\varkappa}=\left.(z{\mathfrak{p}}_{\varkappa}(z))\right|_{z=0}=\begin{cases}\begin{pmatrix}a_{{\varkappa}-1}&0\\ a_{{\varkappa}}&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&a_{{\varkappa}}\\ 0&a_{{\varkappa}+1}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$ (2.10)
Proof.
Let us apply ${\mathcal{M}}_{\varkappa}$ to $g_{z}(x)$. To this end we consider the function
$\frac{1}{2}{\mathfrak{p}}_{{\varkappa}}(z)\frac{AT^{\prime}(z)}{x-T(z)}{\mathfrak{p}}^{*}_{{\varkappa}}(z)=({\mathcal{M}}_{\varkappa}g_{z})(x)+\frac{B_{0}(x)}{z}$
where
$B_{0}(x)=\frac{\Phi_{\varkappa}A\Phi^{*}_{\varkappa}}{x-T(0)}.$
Set $z=T^{\circ k}(0)$ and we obtain (2.9). ∎
Remark 2.5.
Define the Herglotz function $w^{n}_{x}(z)$ in $z$, $n\geq 0$, by
$w^{n}_{x}(z):=\frac{1}{2^{n}}\frac{(T^{\circ n})^{\prime}(z)}{x-T^{\circ n}(z)}=\frac{zT(z)\cdots T^{\circ(n-1)}(z)}{{x-T^{\circ n}(z)}},\quad x\in[-\xi,\xi],$
where by definition $T^{\circ 0}(z)=z$. We can rewrite (2.9) in terms of $w^{k}_{x}(T(0))$ as follows
${\mathcal{M}}_{\varkappa}w_{x}^{k-1}(T(0))A=w_{0}^{k-1}(T(0))w_{x}^{0}(T(0))\Phi_{\varkappa}A\Phi_{\varkappa}^{*}$
$+w_{x}^{k}(T(0)){\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))A{\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))^{*}.$ (2.11)
3. Main theorem
Note that
${\mathsf{E}}\subset[-\xi,\xi],\quad\xi=\frac{1+\sqrt{1+4\lambda}}{2}<\lambda,$
and these are exact upper and lower bounds for ${\mathsf{E}}$.
Lemma 3.1.
Let ${\varkappa}^{\prime}_{n}$ be a certain “extension” of ${\varkappa}$, that is, ${\varkappa}_{n}^{\prime}$ is such that
$({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}^{*}\Sigma_{{\varkappa}_{n}^{\prime}})(dx)=\Sigma_{{\varkappa}}(dx),\quad{\mathfrak{M}}_{{\varkappa}^{\prime}_{n}}={\mathcal{M}}_{\hat{\underline{s}}^{n-1}{\varkappa}_{n}^{\prime}}\cdots{\mathcal{M}}_{{\varkappa}_{n}^{\prime}}.$
Define
$h_{n}(x)=({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x),\quad h_{0}(x)=w_{x}^{0}(T(0))I.$
Then the sequence $h_{n}(x)$ is uniformly bounded on $[-\xi,\xi]$.
Proof.
Acting one by one by the operators ${\mathcal{M}}_{\hat{\underline{s}}^{k}{\varkappa}_{n}^{\prime}}$ and applying (2.5) we obtain that $h_{n}(x)$ is of the form
$h_{n}(x)=\sum_{k=0}^{n}A_{k}^{n}w^{k}_{x}(T(0)),$
where $A_{k}^{n}$ are non-negative matrix coefficients.
Further, on $[-\xi,\xi]$ the function
$w^{0}_{x}(T(0))=\frac{1}{x-T(0)}=\frac{1}{x+\lambda}$
is two-sided bounded
$\frac{1}{\lambda+\xi}\leq w^{0}_{x}(T(0))\leq\frac{1}{\lambda-\xi}.$
Since
$w^{k}_{x}(T(0))=\frac{1}{2^{k}}\frac{(T^{\circ k})^{\prime}(T(0))}{x-T^{\circ k}(T(0))}=\frac{1}{2^{k}}\sum_{T^{\circ k}(y)=x}\frac{1}{y-T(0)}=\frac{1}{2^{k}}\sum_{T^{\circ k}(y)=x}\frac{1}{y+\lambda}$
and $\{y:T^{\circ k}(y)=x\}\subset[-\xi,\xi]$, we have
$\frac{1}{\lambda+\xi}\leq w^{k}_{x}(T(0))\leq\frac{1}{\lambda-\xi}$ (3.1)
for all $k\geq 0$.
Using the upper bound in (3.1) we have
$\text{\rm tr}\,h_{n}(x)\leq\frac{1}{\lambda-\xi}\sum_{k=0}^{n}\text{\rm tr}A_{k}^{n}.$ (3.2)
On the other hand by definition
$\int\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int\text{\rm tr}(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))$
and therefore
$\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))=\int\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))\geq\frac{1}{\lambda+\xi}\sum_{k=0}^{n}\text{\rm tr}(A^{n}_{k}\int\Sigma_{{\varkappa}_{n}^{\prime}}(dx))$
Since all matrix measures are normalized
$\int\Sigma_{{\varkappa}_{n}^{\prime}}(dx)=I,$
we obtain
$\sum_{k=0}^{n}\text{\rm tr}A_{k}^{n}\leq(\lambda+\xi)\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx)).$ (3.3)
Combining (3.2) and (3.3) we get
$\text{\rm tr}\,h_{n}(x)\leq\frac{\lambda+\xi}{\lambda-\xi}\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))$ (3.4)
Proof of the Main Theorem 1.1.
Assume that for some ${\varkappa}\in{\mathbb{Z}}_{2}$ and $x_{0}\in{\mathsf{E}}$
$\text{\rm tr}\Sigma_{{\varkappa}}(\{x_{0}\})>0.$ (3.5)
Note that for a test function $g(x)$
$\int g(x)\text{\rm tr}\left((h_{n}(x)\Sigma_{{\varkappa}^{\prime}_{n}}(dx)\right)=$ $\int g(x)\text{\rm tr}\left(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x)\Sigma_{{\varkappa}^{\prime}_{n}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0}\cdot g\circ T^{\circ n})(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(h_{0}(x)g(T^{\circ n}(x))({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}^{*}\Sigma_{{\varkappa}_{n}^{\prime}})(dx)\right)$
$=$ $\int g(T^{\circ n}(x))\text{\rm tr}\left(h_{0}(x)\Sigma_{\varkappa}(dx)\right)$ (3.6)
We choose a subsequence $I_{x_{0}}$ such that
$\lim_{\begin{subarray}{c}n\to\infty\\ n\in I_{x_{0}}\end{subarray}}T^{\circ n}(x_{0})=y_{0}.$
Now, let us fix an arbitrary small ${\epsilon}$-vicinity of $y_{0}$, $V_{\epsilon}=({y_{0}-{\epsilon}},{y_{0}+{\epsilon}})$. For a sufficiently big $n\in I_{x_{0}}$ we have $T^{\circ n}(x_{0})\in V_{\epsilon}$ and threrefore
$\int_{V_{\epsilon}}\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int_{(T^{\circ n})^{-1}V_{\epsilon}}\text{\rm tr}(h_{0}(x)\Sigma_{{\varkappa}}(dx))\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
In a combination with (3.4) we have
$C\text{\rm tr}\Sigma_{{\varkappa}_{n}^{\prime}}(V_{\epsilon})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
with a uniform constant $C$.
Let ${\varkappa}_{n}^{\prime}=2^{n}{\varkappa}$. Passing to the limit in $n$ we get ${\varkappa}_{n}^{\prime}\to 0$ and
$C\text{\rm tr}\Sigma_{0}(V_{\epsilon})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\})).$
Since ${\epsilon}$ is arbitrary small, the limit measure $\Sigma_{0}$ associated to the initial $J_{0}$ has a non-trivial mass at $y_{0}$. But the spectrum of $J_{0}$ is singular continuous. We arrive to a contradiction, that is, $\Sigma_{{\varkappa}}(\{x_{0}\})=0$. ∎
Appendix A Some explicit formulas
A.1. Recursion for resolvent matrix functions
Lemma A.1.
Let
$W_{\varkappa}(z)=\int\frac{\Sigma_{\varkappa}(dx)}{x-z}$
be the resolvent matrix-function. The following recursion holds
$W_{\varkappa}(z)=-\frac{1}{z}\pi_{\varkappa}+z{\mathfrak{p}}_{\varkappa}(\bar{z})^{*}W_{\hat{}\underline{s}{\varkappa}}(T(z)){\mathfrak{p}}_{\varkappa}(z)$ (A.1)
where
$\pi_{\varkappa}=\begin{cases}\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&0\\ 0&1\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$
Proof.
Recall that matrix orthogonal polynomials of the second kind are given by
$\int\frac{\Sigma(dx)}{x-z}\begin{pmatrix}p_{n}^{-}(x)-p_{n}^{-}(z)\\ -a_{0}(q_{n}^{-}(x)-q_{n}^{-}(z))\end{pmatrix}=\begin{pmatrix}q_{n}^{-}(z)\\ 0\end{pmatrix},$ (A.2)
$\int\frac{\Sigma(dx)}{x-z}\begin{pmatrix}-a_{0}(q_{n}^{+}(x)-q_{n}^{+}(z))\\ p_{n}^{+}(x)-p_{n}^{+}(z)\end{pmatrix}=\begin{pmatrix}0\\ q_{n}^{+}(z)\end{pmatrix}.$ (A.3)
According to this we define
${\mathfrak{q}}_{\varkappa}=\begin{cases}\begin{pmatrix}\frac{1}{a_{{\varkappa}-1}}&0\\ 0&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&0\\ 0&\frac{1}{a_{{\varkappa}+1}}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$
Starting with (2.4), using orthogonality and (A.2)–(A.3), we have
$\frac{1}{2}\int\frac{T^{\prime}(z)}{x-T(z)}\Sigma_{\hat{}\underline{s}{\varkappa}}(dx)=\int\frac{{\mathfrak{p}}^{-1}_{\varkappa}(x)^{*}\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}}{x-z}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}}{x-z}+\int\frac{({\mathfrak{p}}^{-1}_{\varkappa}(x)^{*}-{\mathfrak{p}}^{-1}_{\varkappa}(z)^{*})}{x-z}\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)}{x-z}{\mathfrak{p}}^{-1}_{\varkappa}(z)+{\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)({\mathfrak{p}}_{\varkappa}(x)^{-1}-{\mathfrak{p}}_{\varkappa}(z)^{-1})}{x-z}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)}{x-z}{\mathfrak{p}}^{-1}_{\varkappa}(z)+{\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}{\mathfrak{q}}_{\varkappa}.$
We get (A.1). ∎
Lemma A.2.
Let ${\mathfrak{P}}^{n}_{\varkappa}(z)$ denote the product of matrix-functions
${\mathfrak{P}}^{n}_{\varkappa}(z)={\mathfrak{p}}_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ n}(z))\cdots{\mathfrak{p}}_{\hat{}\underline{s}{\varkappa}}(T(z)){\mathfrak{p}}_{{\varkappa}}(z)$
Then
$1=w_{0}^{0}(T(0)){\langle\Psi_{\varkappa},\Upsilon_{\hat{\underline{s}}{\varkappa}}\rangle^{2}}+\sum_{k=2}^{\infty}w^{k-1}_{0}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{\varkappa},\Upsilon_{\hat{}\underline{s}^{k}{\varkappa}}\rangle^{2}$ (A.4)
where
$\Psi_{\varkappa}=\begin{cases}\begin{pmatrix}a_{{\varkappa}-1}\\ a_{{\varkappa}}\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}a_{{\varkappa}}\\ a_{{\varkappa}+1}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}\quad\Upsilon_{\varkappa}^{*}=\begin{cases}\begin{pmatrix}1&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&1\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$
Proof.
Iterating (A.1) we get
$W_{\varkappa}(z)=$ $-\frac{1}{z}\pi_{\varkappa}+\sum_{k=1}^{n-1}w^{k}_{0}(z){\mathfrak{P}}^{k-1}_{\varkappa}(\bar{z})^{*}\pi_{\hat{}\underline{s}^{k}{\varkappa}}{\mathfrak{P}}^{k-1}_{\varkappa}(z)$
$+$ $\frac{(T^{\circ n})^{\prime}(z)}{2^{n}}({\mathfrak{P}}^{n-1}_{\varkappa}(\bar{z}))^{*}W_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ n}(z)){\mathfrak{P}}^{n-1}_{\varkappa}(z)$ (A.5)
Since $W_{\varkappa}(z)$ is regular at the origin we get the following relation
$\pi_{\varkappa}=$ $\frac{1}{-T(0)}\Phi_{\varkappa}^{*}\pi_{\hat{\underline{s}}{\varkappa}}\Phi_{\varkappa}+\Phi_{\varkappa}^{*}\sum_{k=2}^{n-1}w^{k-1}_{0}(T(0)){\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}\pi_{\hat{}\underline{s}^{k}{\varkappa}}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Phi_{\varkappa}$
$+$ $\Phi_{\varkappa}^{*}{\mathfrak{P}}^{n-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}\int w^{n-1}_{x}(T(0))\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx){\mathfrak{P}}^{n-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Phi_{\varkappa}$
where
$\Phi_{\varkappa}=\left.(z{\mathfrak{p}}_{\varkappa}(z))\right|_{z=0}=\Psi_{\varkappa}\Upsilon_{\varkappa}^{*}.$ (A.6)
Note that $\Upsilon_{\varkappa}\Upsilon_{\varkappa}^{*}=\pi_{\varkappa}$ and all summand in the RHS are positive. ∎
A.2. Periodic case
In this section we derive an explicit formula for eigenfunction of a matrix Ruelle operator if ${\varkappa}$ is periodic, i.e., there exists $d$ such that $\hat{\underline{s}}^{d}{\varkappa}={\varkappa}$.
The computations are based essentially on the identity (A.4). The measure $\Sigma_{\varkappa}(dx)$ is the eigen-measure for the adjoint to the Ruelle operator
$({\mathcal{M}}_{\varkappa}^{*}\cdots{\mathcal{M}}^{*}_{\hat{\underline{s}}^{d-1}{\varkappa}}\Sigma_{{\varkappa}})(dx)=\Sigma_{\varkappa}(dx).$
Of course the same is true for all shifts $\Sigma_{\hat{\underline{s}}^{k}{\varkappa}}(dx)$, $k=1,...,d-1$.
Having in mind Lemma 2.4, we are looking for the eigen-matrix-function $h_{\varkappa}(x)$,
$({\mathcal{M}}_{\hat{\underline{s}}^{d-1}{\varkappa}}\cdots{\mathcal{M}}_{\varkappa}h_{\varkappa})(x)=h_{\varkappa}(x),$ (A.7)
in the form
$h_{\varkappa}(x)=\frac{A_{1}}{x-T(0)}+\sum_{k=2}^{\infty}w_{x}^{k-1}(T(0))A_{k}.$
Due to (2.9),
$({\mathcal{M}}_{\varkappa}h_{\varkappa})(x)=$ $\frac{\Phi_{\varkappa}A_{1}\Phi_{{\varkappa}}^{*}}{-T(0)(x-T(0))}+\sum^{\infty}_{k=2}w_{0}^{k-1}(T(0))\frac{\Phi_{\varkappa}A_{k}\Phi_{{\varkappa}}^{*}}{x-T(0)}$
$+$ $\sum_{k=2}^{\infty}w_{x}^{k-1}(T(0)){\mathfrak{p}}_{\varkappa}(T^{\circ(k-1)}(0))A_{k-1}{\mathfrak{p}}_{\varkappa}(T^{\circ(k-1)}(0))^{*}.$
Similarly, applying ${\mathcal{M}}_{\hat{\underline{s}}{\varkappa}},\dots,{\mathcal{M}}_{\hat{\underline{s}}^{d-1}{\varkappa}}$, we obtain that (A.7) yields the recursion
$A_{k}={\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(k-1)}(0))\cdots{\mathfrak{p}}_{\varkappa}(T^{\circ(k-d)}(0))A_{k-d}$
×$\times\left({\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(k-1)}(0))\cdots{\mathfrak{p}}_{\varkappa}(T^{\circ(k-d)}(0))\right)^{*},\quad\text{for}\ k\geq d+1.$ (A.8)
Now, from first $d$ equations we can observe that $A_{1},\dots,A_{d}$ are necessarily of the form
$A_{1}=y_{1}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}^{*},\quad A_{2}=y_{2}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}^{*}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))^{*},$
and so on… till
$A_{d}=y_{d}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(d-1)}(0))\cdots{\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}\Psi_{{\varkappa}}^{*}$
×$\times\left({\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(d-1)}(0))\cdots{\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\right)^{*}$ (A.9)
with scalars $y_{1},\dots,y_{d}$. In combination with (A.2) we obtain, $l\geq 1$,
  • (1)
    for $k=ld+1$
    $A_{k}=y_{1}{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))^{*},$
  • (2)
    for $k=ld+2$
    $A_{k}=y_{2}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))^{*},$
  • (d)
    and so on… till $k=ld+d$
    $A_{k}=y_{d}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}\Psi_{{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}.$
In this way we reduce (A.7) to a finite linear system
$\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}=L\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}=\begin{pmatrix}L_{11}&\ldots&L_{1d}\\ \vdots&&\vdots\\ L_{d1}&\ldots&L_{dd}\end{pmatrix}\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}$
with the matrix entries: for $j\not=d$
$L_{jn}=\sum_{k=ld+n-j}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2},\ n\not=j+1,$
$L_{j(j+1)}=\frac{\langle\Psi_{\hat{\underline{s}}^{d-j-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=ld+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-j}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-j-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2},$
and
$L_{dn}=\sum_{k=ld+n}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2},\quad n\geq 2,$
$L_{d1}=\frac{\langle\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2}}{-T(0)}+\sum_{k=dl+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2}.$
For a fixed $n$ we have
$\sum_{j=1}^{n}L_{jn}=$
$\frac{\langle\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=ld+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}$
$+\sum_{\begin{subarray}{c}k=ld+n-j\\ n\not=j+1\end{subarray}}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2}$
$=\frac{\langle\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=2}^{\infty}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+k}{\varkappa}}\rangle^{2}$
By (A.4) we obtain
$\sum_{j=1}^{n}L_{jk}=1.$
In other words we get a row-eigenvector $e$ of the matrix $L$,
$eL=e,\quad e=\begin{pmatrix}1&\ldots&1\end{pmatrix}.$
By the classical Perron-Frobenius Theorem $1$ is the spectral radius of the matrix $L$ and there is a unique column-eigenvector $f$, $Lf=f$, with positive entries $\{y_{j}\}_{1}^{d}$ normalized by the condition
$ef=\sum_{j=1}^{d}y_{j}=1.$
Thus we obtain the eigen-matrix-function $h_{\varkappa}(x)$, see (A.7), (A.2) and (1)–(d).
Remark A.3.
As a result we get an invariant measure,
$\mu(dx)=\text{\rm tr}(h_{\varkappa}(x)\Sigma_{\varkappa}(dx)),\quad\int g(T(x))\mu(dx)=\int g(x)\mu(dx).$
Thus we have a “computational” proof of the main theorem in the given particular case. It is easy to see that such a measure $\mu$ has no mass-points. Indeed, assume that $\mu(\{x_{0}\})=\delta>0$. Thus
$\mu(\{x_{0}\})=\sum_{T(y)=x_{0}}\mu(\{y\})$
If $x_{0}$ is not a fixed or periodic point preimages of different generations do not intersect, and we get that the total mass of them is infinite. If $x_{0}$ is periodic, then among preimages we have non-periodic $y_{0}$ and we apply the same arguments to this point. Note that the vectors
${\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}=\begin{pmatrix}1&a_{\hat{\underline{s}}{\varkappa}}\\ 0&a_{\hat{\underline{s}}{\varkappa}+1}\end{pmatrix}\begin{pmatrix}1&0\\ 0&1/{T(0)}\end{pmatrix}\begin{pmatrix}a_{{\varkappa}-1}\\ a_{{\varkappa}}\end{pmatrix},\quad\Psi_{\hat{\underline{s}}{\varkappa}}=\begin{pmatrix}a_{\hat{\underline{s}}{\varkappa}}\\ a_{\hat{\underline{s}}{\varkappa}+1}\end{pmatrix}$
are linearly independent, therefore $h_{\varkappa}(x)>0$. Since $\mu(dx)$ does not contain mass-point components and $h_{\varkappa}(x)>0$ we get that $\Sigma_{\varkappa}(dx)$ also has no mass-point components.
A.3. Computations for $h_{k}(x)$ in a general case
We introduce $A\star B=ABA^{*}$. Then, see (A.2),
${\mathcal{M}}_{2^{n-k}{\varkappa}}\cdots{\mathcal{M}}_{2^{n-1}{\varkappa}}{\mathcal{M}}_{2^{n}{\varkappa}}w_{x}^{0}(T(0))I=w_{x}^{0}(T(0))f^{n-k}_{0}\Phi_{2^{n-k}{\varkappa}}\star I$
$+w_{x}^{1}(T(0))f^{n-k}_{1}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T(0))\Phi_{2^{n-k+1}{\varkappa}})\star I+\dots$
$+w_{x}^{j}(T(0))f^{n-k}_{j}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})\star I$
$+\dots+w_{x}^{k}(T(0))f^{n-k}_{k}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ k}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T(0))\Phi_{2^{n}{\varkappa}})\star I$
$+w_{x}^{k+1}(T(0))({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0)){\mathfrak{p}}_{2^{n-1}{\varkappa}}(T^{\circ 2}(0)){\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))\star I$
with certain constants $f^{n-k}_{j}$, $0\leq j\leq k$. Since, due to (2.5),
${\mathcal{M}}_{2^{n-(k+1)}{\varkappa}}w_{x}^{j}(T(0))A=$ $w_{0}^{j}(T(0))w_{x}^{0}(T(0))\Phi_{2^{n-(k+1)}{\varkappa}}\star A$
$+$ $w_{x}^{j+1}(T(0)){\mathfrak{p}}_{2^{n-(k+1)}{\varkappa}}(T^{\circ(j+1)}(0))\star A,$
these constants obey the following recurrence relations:
$f_{j+1}^{n-(k+1)}=f_{j}^{n-k}$
and
$f_{0}^{n-(k+1)}=w_{0}^{k+1}(T(0))\|({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))^{*}\Upsilon\|^{2}$
$+\sum_{j=0}^{k}w_{0}^{j}(T(0))\|({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})^{*}\Upsilon\|^{2}f_{j}^{n-k}.$
Since
$({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})^{*}\Upsilon=\frac{a_{2^{n-k}{\varkappa}-1}\cdots a_{2^{n-k+j}{\varkappa}-1}}{T^{j}(0)\cdots T(0)}\Upsilon$
and
$({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))^{*}\Upsilon=\frac{a_{2^{n-k}{\varkappa}-1}\cdots a_{2^{n}{\varkappa}-1}}{T^{k+1}(0)\cdots T(0)}\Upsilon$
we obtain
$f_{0}^{n-(k+1)}=$ $w_{0}^{k+1}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{k+1}(0)\cdots T(0))^{2}}$
$+$ $\sum_{j=0}^{k}w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n-k+j}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}f_{j}^{n-k}$
$=$ $w_{0}^{k+1}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{k+1}(0)\cdots T(0))^{2}}$
$+$ $\sum_{j=0}^{k}w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-k}{\varkappa}-1}\cdots a^{2}_{2^{n-k+j}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}f_{0}^{n-k+j}.$
With changes of variables we get
$f_{0}^{n-j}=$ $w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}}$
$+$ $\sum_{s=0}^{j-1}w_{0}^{j-s-1}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n-s}{\varkappa}-1}}{(T^{j-s-1}(0)\cdots T(0))^{2}}f_{0}^{n-s}.$ (A.10)
We define vectors
${\mathfrak{f}}^{n}_{k}=\begin{pmatrix}f_{0}^{n}\\ \vdots\\ f^{n-k}_{0}\end{pmatrix},\quad{\mathfrak{g}}^{n}_{k}=\begin{pmatrix}g_{0}^{n}\\ \vdots\\ g^{n-k}_{0}\end{pmatrix},$
where
$g_{0}^{n}=f_{0}^{n}=w_{0}^{0}(T(0)),\quad g_{0}^{n-j}=w_{0}^{j}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n}{\varkappa}-1}}{(T^{j}(0)\cdots T(0))^{2}},\quad 1\leq j\leq k.$
Also we define a lower-triangular matrix ${\mathfrak{L}}^{n}_{k}$ with the entries
$L_{js}=w_{0}^{j-s-1}(T(0))\frac{a^{2}_{2^{n-j+1}{\varkappa}-1}\cdots a^{2}_{2^{n-s}{\varkappa}-1}}{(T^{j-s-1}(0)\cdots T(0))^{2}},\quad 0\leq s\leq j-1.$
Then, due to (A.3), for the vector ${\mathfrak{f}}_{k}^{n}$ we get the system
$(I-{\mathfrak{L}}_{k}^{n}){\mathfrak{f}}^{n}_{k}={\mathfrak{g}}_{k}^{n}.$
It is evident that the system is solvable
${\mathfrak{f}}^{n}_{k}=(I+{\mathfrak{L}}_{k}^{n}+\dots+({\mathfrak{L}}_{k}^{n})^{k}){\mathfrak{g}}_{k}^{n}.$
Moreover we get $f^{n-j}_{0}>0$ for all $j$. As soon as the system is solved we have
${\mathcal{M}}_{2^{n-k}{\varkappa}}\cdots{\mathcal{M}}_{2^{n-1}{\varkappa}}{\mathcal{M}}_{2^{n}{\varkappa}}w_{x}^{0}(T(0))I=w_{x}^{0}(T(0))f^{n-k}_{0}\Phi_{2^{n-k}{\varkappa}}\star I$
$+w_{x}^{1}(T(0))f^{n-k+1}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T(0))\Phi_{2^{n-k+1}{\varkappa}})\star I+\dots$
$+w_{x}^{j}(T(0))f^{n-k+j}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})\star I$
$+\dots+w_{x}^{k}(T(0))f^{n}_{0}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ k}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T(0))\Phi_{2^{n}{\varkappa}})\star I$
$+w_{x}^{k+1}(T(0))({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T^{\circ 2}(0)){\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))\star I$